MDRS Photogrammetry Scans Completed

The MarsVR team has completed their scans of the MDRS facilities and terrain.  Several Matterport 360 walkthoughs are already available, and the team will be working this summer to create the full photo-realistic VR environment for MDRS crew training, that will be put into use for next field season.

Read More

Call for help on VR project – June 19th – 25th

Hi all,
To support the creation of a VR environment for future MDRS crew training and research, we will be conducting scans and drone operations at the MDRS starting on June 19th for around a week.  We’d like to get help from a past crewmember at the MDRS.  Please let us know if you are interested in helping out with our project – we can work out your travel arrangements.  Please forward this to anybody that might be interested in participating.
More info on our project is at: http://marsvr.io

Thanks,

James Burk

Program Manager, MarsVR

marsvr@marssociety.org

Operations Report – June 4th

SOL: 0

Name of person filing report: Shannon Rupert

Non-nominal systems: Do you believe in magic? Because at this very moment, we have no non-nominal systems! (Knock wood)

Notes on non-nominal systems: Happy dance!

Generator (hours run): Turned off at 8 am, will turn back on at 8 pm

Solar—100 % SOC

Diesel – 50 %

Propane – 58 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) –Not checked

Water (trailer) –450 gallons.

Water (static) – 450 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: not recorded

Toilet tank emptied: yes, and flushed twice by filling and emptying.

ATV’s Used: Yes

Oil Added? no

ATV Fuel Used: n/a

# Hours the ATVs were Used today: 1

Notes on ATVs: Mostly moving around, some training. Need new bolts for new battery for 300.

Deimos rover used: Yes

Hours:134.0

Beginning charge: 100

Ending charge:100

Currently charging: no

Sojourner rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: yes

Hours: 42.3

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: no

Hours: 31.7

Beginning charge: 100

Ending charge: 100

Currently charging: yes (NOTE: Driver’s seat is bent, should be replayed by one from Phobos)

Curiosity rover used: yes

Hours: 39.0

Beginning charge: 97

Ending charge: still being used

Currently charging: No

HabCar used and why, where? Yes, to get water and move things from MDRS to town, also trash

General notes and comments: The Korean Project began last evening as the first of the team arrived. We did some generally cleaning. This morning the transformation began. We have 40 plus people on site. The Hab was given some snazzy upgrades in terms of lower deck tables and upper deck kitchen. An airstream was brought in to serve as their production center and is parked on the opposite side of the GreenHab and tunnel. Four motorhomes, a bathroom trailer, water truck and large generator are now parked in my yard. Tomorrow they are putting in a misting shelter. I am looking forward to this as I can’t use my AC due to their power needs.

Cameras have been installed everywhere. Portable toilets have all been removed from campus and URC sites.

Summary of internet: Nothing to report

Summary of suits and radios: The crew installing small cameras on the new faceplates and understand that if there is any damage they will be charged $50 per faceplate.

Summary of Hab operations: The AC (swamp cooler) has been installed on the upper deck and it was working wonderfully until the crew emptied the water tank. I got it up and running again but not sure if the water is working correctly. I will check later.

Summary of GreenHab operations: 50 plus new plants growing food will be installed in the GreenHab this evening for use by the crew for food. Plants were grown off station for this project,.

Summary of ScienceDome operations: Lab fridge is clean, both fridge and freezer were relocated to the lab table and off the floor. New lamp left by the WMS crews was installed and looks great.

Summary of RAM operations: The tool box on the lower deck was relocated to its new home in the RAM. We have cleaned the RAM for the filming but it is not organized properly.

Summary of health and safety issues: Because so much trash was generated by the URC, we burned it and now have a bit of a mess to clean up. Broken and burned glass and metal needs to be removed and taken to town.

The road is so sand pitted now that we had many people stuck today.

Questions, concerns and requests to Mission Support: Nothing at this time